OpenStudy (anonymous):

Given that f(x) = 3x + 1 and g(x) = the quantity of 4x plus 2 divided by 3 solve for g(f(0)). 2 6 8 9

OpenStudy (anonymous):

f(x) = 3x + 1 g(x) = \[4x + 2 \over 3\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

g(f(0)) ^^

ganeshie8 (ganeshie8):

first find f(0)

OpenStudy (anonymous):

sorry was afk eating lunch

OpenStudy (anonymous):

f(x) = 3x + 1 F(0) = 3(0) + 1 0 = 0 + 1 0 = 0?

ganeshie8 (ganeshie8):

f(x) = 3x + 1 F(0) = 3(0) + 1 = 0 + 1 = 1

ganeshie8 (ganeshie8):

f(0) = 1

ganeshie8 (ganeshie8):

plug that in g(f(0)) g(1)

OpenStudy (anonymous):

g(1)..

ganeshie8 (ganeshie8):

yup, find it

OpenStudy (anonymous):

4x + 2 / 3 4(1) + 2 / 3 4 + 2 = 6 6/3 2

ganeshie8 (ganeshie8):

Yes !!

ganeshie8 (ganeshie8):

good job :)

OpenStudy (anonymous):

thanks man :D